### v1

# Brook’s type upper bounds for coloring parameters of infinite graphs and Kőnig’s Lemma

Abstract.
We prove that Brook’s type theorems for several coloring parameters of infinite graphs, which are true in $\mathsf{ZFC}$, are not provable in $\mathsf{ZF}$ (i.e., the Zermelo–Fraenkel set theory without the Axiom of Choice ($\mathsf{AC}$)). We show that the statement “If $G$ is an infinite connected graph where the maximum degree $\Delta(G)$ is finite, then the total distinguishing number of $G$ is at most $\lceil\sqrt{\Delta(G)}\rceil$” follows from Kőnig’s lemma (any infinite locally finite connected graph has a ray) in $\mathsf{ZF}$. In $\mathsf{ZF}$, we also formulate new conditions for the existence of distinguishing chromatic number, distinguishing chromatic index, total chromatic number, odd chromatic number, and neighbor-distinguishing index in infinite locally finite connected graphs, which are equivalent to Kőnig’s Lemma.
Key words and phrases:
Axiom of Choice, Total colorings, distinguishing colorings, odd colorings, infinite graph
2020 Mathematics Subject Classification:
Primary 03E25; Secondary 05C63, 05C15, 05C69.
1. Introduction
In 1941, Brooks [5] proved that for any connected undirected graph $G$ with maximum degree $\Delta(G)$, the chromatic number $\chi(G)$ of $G$ is at most $\Delta(G)$, unless $G$ is a complete graph or an odd cycle. Brook’s theorem for connected infinite graphs was proved by Behzad and Radjavi [4] in 1976. Since then several authors obtained general upper bounds for different graph coloring parameters of an infinite graph $G$ (cf. [4, 16, 15, 21, 24]).
Fact 1.1.
($\mathsf{ZFC}$) If $G$ is an infinite connected graph where $\Delta(G)$ is finite, then the following holds:
  1. (1)
    (Behzad and Radjavi; [4, Theorem 3]) The chromatic number $\chi(G)$ is at most $\Delta(G)$.
  2. (2)
    (Behzad and Radjavi; [4, Theorem 7]) The chromatic index $\chi^{\prime}(G)$ is at most $\Delta(G)+1$.
  3. (3)
    (Imrich, Klavžar, and Trofimov; [16, Theorem 2.1]) The distinguishing number $D(G)$ is at most $\Delta(G)$.
  4. (4)
    (Lehner, Pilśniak, and Stawiski; [21, Theorem 3]) If $\Delta(G)\geq 3$, then the distinguishing number $D(G)$ is at most $\Delta(G)-1$.
  5. (5)
    (Pilśniak and Stawiski; [24, Theorem 3]) If $\Delta(G)\geq 3$, then the distinguishing index $D^{\prime}(G)$ is at most $\Delta(G)-1$.
  6. (6)
    (Imrich, Kalinowski, Pilśniak, and Shekarriz; [15, Theorem 3]) The distinguishing chromatic number $\chi_{D}(G)$ is at most $12\Delta(G)-$ 1.
  7. (7)
    (Imrich, Kalinowski, Pilśniak, and Shekarriz; [15, Theorem 14]) The distinguishing chromatic index $\chi^{\prime}_{D}(G)$ is at most $\Delta(G)+2$.
Complete definitions of the graph theoretic terminologies will be given in the Appendix.
1.1. Brooks type upper bounds fails in ZF
Graph coloring parameters can be dramatically affected by the inclusion or the exclusion of $\mathsf{AC}$ (cf. [3, 9, 12, 13, 25, 26, 27, 29]). In 1991, Galvin–Komjáth [9] proved that $\mathsf{AC}$ is equivalent to the statement “Any graph has a chromatic number”. In 2003 and 2004, Shelah and Soifer [26, 27] constructed three graphs whose chromatic number is either 2 or 4 in $\mathsf{ZFC}$ and uncountable (if it exists) in some model of $\mathsf{ZF}$ (e.g., in Solovay’s model from [28, Theorem 1]). Herrlich–Tachtsis [12, Proposition 23] proved that no Russell graph has a chromatic number in $\mathsf{ZF}$ [^1] We refer the reader to Herrlich–Tachtsis [12] for the details concerning Russell graph. The above-mentioned results are the chief motivation for us to study the status of the results in Fact 1.1 and other Brook’s type theorems for coloring parameters if $\mathsf{AC}$ is removed.
Results: In this note, we prove the following:
  1. (1)
    (Theorem 3.3) In $\mathsf{ZF}$, Kőnig’s Lemma (a weak form of $\mathsf{AC}$) implies the statement “If $G$ is an infinite connected graph where $\Delta(G)$ is finite, then the total distinguishing number $D^{\prime\prime}(G)$ is at most $\lceil\sqrt{\Delta(G)}\rceil$, the odd chromatic number $\chi_{o}(G)$ is at most $2\Delta(G)$, the total chromatic number $\chi^{\prime\prime}(G)$ is at most $\Delta(G)+10^{26}$”, and the total distinguishing chromatic number $\chi^{\prime\prime}_{D}(G)$ is at most $\Delta(G)+10^{26}+1$”. Consequently, the above statement is true in $\mathsf{ZFC}$.
  2. (2)
    (Theorem 6.1) The statements (1)-(7) in Fact 1.1 and the statement in (1) fail in $\mathsf{ZF}$.
1.2. New equivalents of Kőnig’s lemma
In 1973, Höft–Howard [10] proved that $\mathsf{AC}$ is equivalent to the statement “Any connected graph contains a partial subgraph, which is a tree”. Delhommé–Morillon [7], Banerjee [1, 2], and Banerjee–Molnár–Gopaulsingh [3] continued the research works in Höft–Howard [10] and studied the role of $\mathsf{AC}$ and Kőnig’s Lemma in the existence of spanning subgraphs, maximal independent sets, minimal edge cover, maximal matching, and minimal dominating set. Recently, Banerjee–Molnár–Gopaulsingh [3, Theorem 4.2] and Stawiski [29] proved that Kőnig’s lemma is equivalent to the statement “Any infinite locally finite connected graph has a chromatic number (resp. a chromatic index, a distinguishing number, a distinguishing index, an irreducible proper coloring)”. The techniques in [3, Theorem 4.2] were mainly inspired by the arguments of Herrlich–Tachtsis [12, Proposition 23].
Results: In this note, the first two authors develop the techniques of [3, Theorem 4.2] substantially and investigate the role of Kőnig’s Lemma in the existence of total chromatic number, and some recently investigated coloring parameters like odd chromatic number [6], distinguishing chromatic number [15], distinguishing chromatic index [15], total distinguishing chromatic number [15], and neighbor-distinguishing index [14]. The following table summarizes the new results (Theorems 4.1, 5.1) as well as the published results [3, 7, 29] in this topic. In the table, $\mathcal{P}_{\text{lf,c}}$(property $X$) denotes “Any infinite locally finite connected graph has property $X$”. [^2] We note that Theorems 3.3, 4.1, and 5.1 are a combined effort of the first two authors. Moreover, Theorem 6.1 is due to all the authors.
Known equivalents of Kőnig’s lemma New equivalents of Kőnig’s lemma
$\mathcal{P}_{\text{lf,c}}$(distinguishing number) $\mathcal{P}_{\text{lf,c}}$(distinguishing chromatic number)
$\mathcal{P}_{\text{lf,c}}$(distinguishing index) $\mathcal{P}_{\text{lf,c}}$(distinguishing chromatic index)
$\mathcal{P}_{\text{lf,c}}$(chromatic number) $\mathcal{P}_{\text{lf,c}}$(total chromatic number)
$\mathcal{P}_{\text{lf,c}}$(chromatic index) $\mathcal{P}_{\text{lf,c}}$(total distinguishing chromatic number)
$\mathcal{P}_{\text{lf,c}}$(irreducible proper coloring) $\mathcal{P}_{\text{lf,c}}$(neighbor-distinguishing index)
$\mathcal{P}_{\text{lf,c}}$(spanning tree) $\mathcal{P}_{\text{lf,c}}$(odd chromatic number)
2. Basics
Definition 2.1.
Let $X$ and $Y$ be two sets. We write:
  1. (1)
    $X\preceq Y$, if there is an injection $f:X\rightarrow Y$.
  2. (2)
    $X$ and $Y$ are equipotent if $X\preceq Y$ and $Y\preceq X$, i.e., if there is a bijection $f:X\rightarrow Y$.
  3. (3)
    $X\prec Y$, if $X\preceq Y$ and $X$ is not equipotent with $Y$.
Definition 2.2.
Without $\mathsf{AC}$, a set $m$ is called a cardinal if it is the cardinality $|x|$ of some set $x$, where $|x|$ = $\{y:y\sim x$ and $y$ is of least rank}}\}} where $y\sim x$ means the existence of a bijection $f:y\rightarrow x$ (see [19, Definition 2.2, p. 83] and [17, Section 11.2]).
Definition 2.3.
A topological space $(X,\tau)$ is called compact if for every $U\subseteq\tau$ such that $\bigcup U=X$ there is a finite subset $V\subseteq U$ such that $\bigcup V=X$.
Let $\omega$ be the set of all natural numbers.
Definition 2.4.
A set $X$ is Dedekind-finite if it satisfies the following equivalent conditions (cf. [12, Definition 1]):
  • $\omega\not\preceq X$,333i.e., there is no injection $f:\omega\rightarrow X$.
  • $A\prec X$ for every proper subset $A$ of $X$.
Definition 2.5.
For every family =$\mathcal{B}=\{B_{i}:i\in I\}$ of non-empty sets, $\mathcal{B}$ is said to have a partial choice function if $\mathcal{B}$ has an infinite subfamily $\mathcal{C}$ with a choice function.
Definition 2.6.
(A list of choice forms).
  1. (1)
    $\mathsf{AC_{fin}^{\omega}}$: Any countably infinite family of non-empty finite sets has a choice function. We recall that $\mathsf{AC_{fin}^{\omega}}$ is equivalent to Kőnig’s Lemma as well as the statement “The union of a countable family of finite sets is countable”.
  2. (2)
    $\mathsf{PAC^{\omega}_{fin}}$: Any countably infinite family of non-empty finite sets has a partial choice function.
Definition 2.7.
In model theory, the language of graphs $\mathcal{L}$ consists of a single binary relational symbol $E$ depicting edges, i.e., =$\mathcal{L}=\{E\}$ and a graph is an $\mathcal{L}$-structure $G=(V,E)$ consisting of a non-empty set $V$ of vertices and the edge relation $E$ on $V$. Let $G=(V,E)$ be an $\mathcal{L}$-structure, $\phi(x_{1},...,x_{n})$ be a first-order $\mathcal{L}$-formula, and let $a_{1},...,a_{n}\in V$ for some $n\in\omega\backslash\{0\}$. We write $G\models\phi(a_{1},...,a_{n})$, if the property expressed by $\phi$ is true in $G$ for $a_{1},...,a_{n}$. Let $G_{1}=(V_{G_{1}},E_{G_{1}})$ and $G_{2}=(V_{G_{2}},E_{G_{2}})$ be two $\mathcal{L}$-structures. We recall that if $j:V_{G_{1}}\to V_{G_{2}}$ is an isomorphism, $\varphi(x_{1},...,x_{r})$ is a first-order $\mathcal{L}$-formula on $r$ variables for some $r\in\omega\backslash\{0\}$, and $a_{i}\in V_{G_{1}}$ for each $1\leq i\leq r$, then by induction on the complexity of formulae, one can see that $G_{1}\models\varphi(a_{1},...,a_{r})$ if and only if $G_{2}\models\varphi(j(a_{1}),...,j(a_{r}))$ (cf. [22, Theorem 1.1.10]).
3. Known Results and Results in ZFC
3.1. Known Results
Theorem 3.1.
($\mathsf{ZF}$) The following holds:
  1. (1)
    (Kalinowski, Pilśniak, and Woźniak; [18, Theorem 2.2]) If $G$ is a finite connected graph of order $n\geq 3$, then $D^{\prime\prime}(G)\leq\lceil\sqrt{\Delta(G)}\rceil$.
  2. (2)
    (Caro, Petruševski, and Škrekovski; [6, Theorem 4.9]) If $G$ is a finite connected graph where $\Delta(G)\geq 3$, then $\chi_{o}(G)\leq 2\Delta(G)$ unless $G$ is the cycle graph $C_{5}$.
  3. (3)
    (Molloy and Reed; [23]) If $G$ is a finite graph, then $\chi^{\prime\prime}(G)\leq\Delta(G)+10^{26}$.
  4. (4)
    (Loeb; [20, Theorem 1]) Let $\{X_{i}\}_{i\in I}$ be a family of compact spaces which is indexed by a set $I$ on which there is a well-ordering $\leq$. If $I$ is an infinite set and there is a choice function $F$ on the collection $\{C$ : C is closed, $C\not=\emptyset,C\subset X_{i}$ for some $i\in I\}$, then the product space $\prod_{i\in I}X_{i}$ is compact in the product topology.
By Theorem 3.1(2), it is possible to see that for any graph $G$ (connected or disconnected) where $\Delta(G)\geq 3$, we have $\chi_{o}(G)\leq 2\Delta(G)$. Moreover, if $\Delta(G)\leq 2$, then $\chi_{o}(G)=2\Delta(G)+1$ if and only if $G$ has $C_{5}$ as a connected component, otherwise $\chi_{o}(G)\leq 2\Delta(G)$.
3.2. Results in ZFC
Observation 3.2.
($\mathsf{ZF}$)Any graph $G=(V_{G},E_{G})$ based on a well-ordered set of vertices has a distinguishing chromatic number, a distinguishing chromatic index, a total chromatic number, a total distinguishing number, a total distinguishing chromatic number, a neighbor-distinguishing index, and an odd chromatic number. Moreover, if $G$ is connected, then it has a breadth-first search (BFS)-spanning tree rooted at $v_{0}$ for any $v_{0}\in V_{G}$.
Proof.
If $V_{G}$ is well-orderable, then $E_{G}\subseteq[V_{G}]^{2}$ is well-orderable as well. Thus, we can color each vertex (resp. edge) of $G$ with a unique color. We can use transfinite recursion without invoking any form of choice and modify the BFS algorithm to find a spanning tree rooted at $v_{0}$. ∎
Theorem 3.3.
($\mathsf{ZF}$+ Kőnig’s Lemma) Let $G=(V_{G},E_{G})$ be an infinite connected graph where $\Delta(G)$ is finite. Then the following holds:
  1. (1)
    $D^{\prime\prime}(G)\leq\lceil\sqrt{\Delta(G)}\rceil$,
  2. (2)
    $\chi_{o}(G)\leq 2\Delta(G)$,
  3. (3)
    $\chi^{\prime\prime}(G)\leq\Delta(G)+10^{26}$,
  4. (4)
    $\chi^{\prime\prime}_{D}(G)\leq\Delta(G)+10^{26}+1$.
Proof.
Pick some $r\in V_{G}$. Let $N^{0}(r)=\{r\}$. For each integer $n\geq 1$, define $N^{n}(r)=\{v\in V_{G}:d_{G}(r,v)=n\}$ where “$d_{G}(r,v)=n$” means there are $n$ edges in the shortest path joining $r$ and $v$. Each $N^{n}(r)$ is finite since $\Delta(G)$ is finite, and $V_{G}=\bigcup_{n\in\omega}N^{n}(r)$ since $G$ is connected. By Kőnig’s Lemma, which is equivalent to $\mathsf{AC^{\omega}_{fin}}$, $V_{G}$ is countably infinite (and hence, well-orderable).
(1). Since $G$ is an infinite connected graph, $\Delta(G)\geq 2$. Let $v_{0}\in V_{G}$ be such that $\deg(v_{0})=\Delta(G)$. By Observation 3.2, there is a $BFS$-spanning tree $T$ of $G$ rooted at $v_{0}$ in $\mathsf{ZF}$. Assign color 00 to $v_{0}$. By Kőnig’s Lemma, there is an infinite ray $p=\langle v_{i}:i\in\omega\rangle$ in $T$ starting from $v_{0}$ since $V_{T}=V_{G}$ is countably infinite. We will construct a total distinguishing coloring with $\lceil\sqrt{\Delta(G)}\rceil$ colors. For a vertex $v\in V_{G}$, we denote
$N(v)=N^{1}(v)\quad\text{ and }\quad S(v)=\{(\{v,u\},u):\{v,u\}\in E_{G}\}.$
Pick $z\in N(v_{0})\setminus\{v_{1}\}$ (since selecting an element from a set does not involve any form of choice) and assign the color pair $(1,1)$ to both ($(\{v_{0},v_{1}\},v_{1})$ and ($(\{v_{0},z\},z)$. Since $\deg(v_{0})=\Delta(G)$, and the pair $(1,1)$ is used twice to color the elements of $S(v_{0})$, we can color every pair from $S(v_{0})\setminus\{(\{v_{0},v_{1}\},v_{1}),(\{v_{0},z\},z)\}$ with distinct pairs of colors different from $(1,1)$. Following a similar strategy as in [18, Theorem 2.2] due to Kalinowski, Pilśniak, and Woźniak, we will color $G$ such that $v_{0}$ will be the only vertex colored with 00, and the pair $(1,1)$ appears twice in $S(v_{0})$. Hence $v_{0}$ will be fixed by every color preserving automorphism, and all vertices in $N(v_{0})$ will be fixed except, possibly $v_{1}$ and $z$. In order to distinguish $v_{1}$ and $z$, we color the elements from
{($\{(\{v_{1},u\},u)\in S(v_{1}):u\in N^{2}(v_{0})\}\quad\text{ and }\quad\{(\{z,u\},u)\in S(z):\{z,u\}\in E_{T},u\in N^{2}(v_{0})\}$
with two distinct sets of color pairs. This is possible since each of these sets contains at most $\Delta(G)-1$ pairs, while at least $\Delta(G)$-many distinct color pairs are available. Now suppose that for some $k\in\omega$, every element from $u\in\bigcup_{i\leq k}N^{i}(v_{0})$ are colored such that they are fixed by every color preserving automorphism. For every $u\in N^{k}(v_{0})$, color ($(\{u,z\},z)$ where $z$ is a descendent of $u$ in $T$, with distinct color pairs except $(1,1)$. Hence all neighbors in $N(u)\cap N^{k+1}(v_{0})$ will be fixed. Therefore the elements of $\bigcup_{i\leq k+1}N^{i}(v_{0})$ are also fixed by every color preserving automorphism.
Continuing in this fashion by mathematical induction (and noting that the process will not stop at a finite stage), we end up coloring the vertices and edges of $T$. Finally, we color the edges in $E_{G}\setminus E_{T}$ with 00. The color assignment of $G$, defined above, is total distinguishing.
(2). By Theorem 3.1(2), every finite subgraph of $G$ is $(2\Delta(G)+1)$-odd colorable. Endow the set $\{1,...,2\Delta(G)+1\}$ with the discrete topology and $T=\{1,...,2\Delta(G)+1\}^{V_{G}}$ with the product topology. Since $V_{G}$ is well-orderable, $\{1,...,2\Delta(G)+1\}\times{V_{G}}$ is well-orderable in $\mathsf{ZF}$. By Theorem 3.1(4), $T$ is compact. We can use Theorem 3.1(2) and the compactness of $T$ to prove that $G$ is $(2\Delta(G)+1)$-odd colorable similarly to the proof of de Bruijn–Erdős theorem, without invoking any form of choice. Now, if $G$ were not $2\Delta(G)$-colorable, then there would exist a finite subgraph $H$ of $G$ with $\chi_{o}(H)=2\Delta(G)+1$. By Theorem 3.1(2), $H$ has $C_{5}$ as a connected component. Then $G$ would have to contain a vertex of degree greater than $\Delta(G)$, which is impossible (see the paragraph after Theorem 3.1).
(3). This follows from Theorem 3.1(3), and a straightforward compactness argument as in (2).
(4). This follows from (3) and the fact that the statement “$\chi^{\prime\prime}_{D}(G)$ is at most $\chi^{\prime\prime}(G)+1$” holds in $\mathsf{ZF}$ since $V_{G}$ is well-orderable (see [15, proof of Theorem 13]). ∎
4. Total, Neighbor-distinguishing, and distinguishing proper colorings
In this section, we prove the following.
Theorem 4.1.
($\mathsf{ZF}$)The following statements are equivalent:
  1. (1)
    Kőnig’s Lemma.
  2. (2)
    Any infinite locally finite connected graph has a distinguishing chromatic number.
  3. (3)
    Any infinite locally finite connected graph has a distinguishing chromatic index.
  4. (4)
    Any infinite locally finite connected graph has a total chromatic number.
  5. (5)
    Any infinite locally finite connected graph has a total distinguishing chromatic number.
  6. (6)
    Any infinite locally finite connected graph has a neighbor-distinguishing index.
Let $G=(V_{G},E_{G})$ be an infinite locally finite connected graph. Then $V_{G}$ is countably infinite (and hence, well-orderable) by Kőnig’s Lemma (see the first paragraph of the proof of Theorem 3.3). Thus, (1) implies (2)-(6) by Observation 3.2. We show that each of the condition (2)-(6) implies (1). Since $\mathsf{AC_{fin}^{\omega}}$ is equivalent to its partial version $\mathsf{PAC_{fin}^{\omega}}$ (Every countably infinite family of non-empty finite sets has an infinite subfamily with a choice function) (cf. [11], [8, the proof of Theorem 4.1(i)]), it suffices to show that (2)-(6) implies $\mathsf{PAC_{fin}^{\omega}}$.
Let $\mathcal{A}=\{A_{n}:n\in\omega\}$ be a countably infinite set of non-empty finite sets without a partial choice function. Without loss of generality, we assume that $\mathcal{A}$ is disjoint. It is enough to show that there exists an infinite locally finite connected graph $G$ such that $G$ does not have the coloring parameters mentioned in (2)-(6).
4.1. Construction of the graph
Pick a countably infinite sequence $T=\{t_{n}:n\in\omega\}$ disjoint from $A=\bigcup_{i\in\omega}A_{i}$. Pick $t^{\prime},t^{\prime\prime}\not\in(\bigcup_{n\in\omega}A_{n})\cup T$ such that $t\neq t^{\prime}$ and consider the following infinite locally finite connected graph $G=(V_{G},E_{G})$ (see Figure 1):
$V_{G}$ $:=(\bigcup_{n\in\omega}A_{n})\cup T\cup\{t^{\prime},t^{\prime\prime}\},$
$E_{G}$ :=$:=\{t^{\prime\prime},t^{\prime}\}\cup\{t^{\prime},t_{0}\}\cup\big{\{}\{t_{n},t_{n+1}\}:n\in\omega\big{\}}$$\cup\big{\{}\{t_{n},x\}:n\in\omega,x\in A_{n}\big{\}}.$
[…]
Figure 1. Graph $G$, an infinite locally finite connected graph.
4.2. G has no distinguishing chromatic number
Claim 4.2.
Any automorphism fixes the vertices $t^{\prime},t^{\prime\prime}$, and $t_{m}$ for each non-negative integer $m$.
Proof.
Fix non-negative integers $n,m,r$. The first-order $\mathcal{L}$-formula
$\mathsf{Deg}_{n}(x):=\exists x_{0}\dots\exists x_{n-1}\big{(}\bigwedge_{i\neq j}^{n-1}x_{i}\neq x_{j}\wedge\bigwedge_{i<n}x\neq x_{i}\wedge\bigwedge_{i<n}Exx_{i}\wedge\forall y(Exy\to\bigvee_{i<n}y=x_{i})\big{)}$
expresses the property that a vertex $x$ has degree $n$, where $Eab$ denotes the existence of an edge between vertices $a$ and $b$. We define the following first-order $\mathcal{L}$-formula:
$\varphi(x):=\mathsf{Deg}_{1}(x)\wedge\exists y(Exy\wedge\mathsf{Deg}_{2}(y))$.
We can see the following:
  1. (i)
    $t^{\prime\prime}$ is the unique vertex such that $G\models\varphi(t^{\prime\prime})$. This means $t^{\prime\prime}$ is the unique vertex such that $\deg(t^{\prime\prime})=1$ and $t^{\prime\prime}$ has a neighbor of degree $2$.
  2. (ii)
    $t^{\prime}$ is the unique vertex such that $G\models\mathsf{Deg}_{2}(t^{\prime})$. So $t^{\prime}$ is the unique vertex with $\deg(t^{\prime})=2$.
Fix any automorphism $\tau$. Since any automorphism preserves the properties mentioned in (i)–(ii), the vertices $t^{\prime}$ and $t^{\prime\prime}$ are fixed by $\tau$. The vertices $t_{m}$ are fixed by $\tau$ by induction as follows: Since $t_{i}$ is the unique vertex of path length $i+1$ from $t^{\prime\prime}$ such that the degree of $t_{i}$ is greater than 1, where $i\in\{0,1\}$, we have that $t_{0}$ and $t_{1}$ are fixed by $\tau$. Assume that $\tau(t_{l})=t_{l}$ for all $l<m-1$. We show that $\tau(t_{m})=t_{m}$. Now, $\tau(t_{m})$ is a neighbor of $\tau(t_{m-1})=t_{m-1}$ which is of degree at least $2$, so $\tau(t_{m})$ must be either $t_{m-2}$ or $t_{m}$, but $t_{m-2}=\tau(t_{m-2})$ is already taken. So, $\tau(t_{m})=t_{m}$. ∎
Claim 4.3.
Fix $p\in\omega$ and $x\in A_{p}$. Then $Orb_{Aut(G)}(x)=\{g(x):g\in Aut(G)\}=A_{p}$.
Proof.
This follows from Claim 4.2 and the fact that each $y\in\bigcup_{n\in\omega}A_{n}$ has path length $1$ from $t_{p}$ if and only if $y\in A_{p}$. ∎
Claim 4.4.
If $f:V_{G}\rightarrow C$ is a distinguishing proper vertex coloring of $G$, then the following holds:
  1. (1)
    If $x,y\in A_{i}$ such that $x\neq y$ for any $i\in\omega$, then $f(x)\neq f(y)$.
  2. (2)
    For each $c\in C$, the set $M_{c}=f^{-1}(c)\cap\bigcup_{i\in\omega}A_{i}$ must be finite.
  3. (3)
    $f[\bigcup_{n\in\omega}A_{n}]$ is infinite.
  4. (4)
    $f[\bigcup_{n\in\omega}A_{n}]$ is Dedekind-finite.
  5. (5)
    For each natural number $m$, if $c_{0},...,c_{m}\in f[\bigcup_{n\in\omega}A_{n}]$, then there is a vertex coloring $h:\bigcup_{n\in\omega}A_{n}\to f[\bigcup_{n\in\omega}A_{n}]\setminus\{c_{0},...,c_{m}\}$ such that $h(x)=f(x)$ whenever $f(x)\neq c_{i}$.
Proof.
(1). This follows by Claim 4.3.
(2). If $M_{c}$ is infinite, then it will generate a partial choice function for $\mathcal{A}$ by (1).
(3). If $f[\bigcup_{n\in\omega}A_{n}]$ is finite, then $\bigcup_{n\in\omega}A_{n}=\bigcup_{c\in f[\bigcup_{n\in\omega}A_{n}]}M_{c}$ is finite by (2), since the finite union of finite sets is finite in $\mathsf{ZF}$. Thus, we obtain a contradiction.
(4). First, we note that $\bigcup_{n\in\omega}A_{n}$ is Dedekind-finite since $\mathcal{A}$ has no partial choice function. For the sake of contradiction, assume that $C=\{c_{i}:i\in\omega\}$ is a countably infinite subset of $f[\bigcup_{n\in\omega}A_{n}]$. Fix a well-ordering $<$ of $\mathcal{A}$ (since $\mathcal{A}$ is countable, and hence well-orderable). Define $d_{i}$ to be the unique element of $M_{c_{i}}\cap A_{n}$ where $n$ is the $<$-least element of $\{m\in\omega:M_{c_{i}}\cap A_{m}\neq\emptyset\}$. Such an $n$ exists since $c_{i}\in f[\bigcup_{n<\omega}A_{n}]$ and $M_{c_{i}}\cap A_{n}$ has a single element by (1). Then $\{d_{i}:i\in\omega\}$ is a countably infinite subset of $\bigcup_{n\in\omega}A_{n}$ which contradicts the fact that $\bigcup_{n\in\omega}A_{n}$ is Dedekind-finite.
(5). Fix some $c_{0}\in f[\bigcup_{n\in\omega}A_{n}]$. Then $Index(M_{c_{0}})=\{n\in\omega:M_{c_{0}}\cap A_{n}\neq\emptyset\}$ is finite. By (3), there exists some $b_{0}\in(f[\bigcup_{n\in\omega}A_{n}]\backslash\bigcup_{m\in Index(M_{c_{0}})}f[A_{m}])$ since the finite union of finite sets is finite (see Figure 2).
[…]
Figure 2. There exists a color $b_{0}$ from the set $(f[\bigcup_{n\in\omega}A_{n}]\backslash\bigcup_{m\in Index(M_{c_{0}})}f[A_{m}])$.
Define a vertex coloring $g:\bigcup_{n\in\omega}A_{n}\rightarrow(f[\bigcup_{n\in\omega}A_{n}]\backslash\{c_{0}\})$ as follows:
$g(x)=\begin{cases}f(x)&\text{if}\,f(x)\neq c_{0},\\ b_{0}&\text{otherwise}.\end{cases}$
Similar to the definition of $g$, if $c_{0},...,c_{m}\in f[\bigcup_{n\in\omega}A_{n}]$, then one can define a vertex coloring $h:\bigcup_{n\in\omega}A_{n}\rightarrow(f[\bigcup_{n\in\omega}A_{n}]\backslash\{c_{0},...,c_{m}\})$ saitisfying the properties of the assertion. ∎
Claim 4.5.
$G$ has no distinguishing chromatic number.
Proof.
Assume that $G$ has a distinguishing chromatic number (see Appendix). Let $f:V_{G}\to C$ be a distinguishing proper vertex coloring with $|C|=\kappa$, where $\kappa$ is the distinguishing chromatic number of $G$. For some $c_{0},c_{1},c_{2}\in f[\bigcup_{n\in\omega}A_{n}]$, we can define a vertex coloring $h:\bigcup_{n\in\omega}A_{n}\rightarrow(f[\bigcup_{n\in\omega}A_{n}]\backslash\{c_{0},c_{1},c_{2}\})$ as in Claim 4.4(5). Let $h(t^{\prime\prime})=c_{0}$, $h(t^{\prime})=c_{1}$, $h(t_{2n})=c_{0}$, and $h(t_{2n+1})=c_{1}$ for all $n\in\omega$. By Claim 4.2, $h:V_{G}\rightarrow(f[\bigcup_{n\in\omega}A_{n}]\backslash\{c_{2}\})$ is a $f[\bigcup_{n\in\omega}A_{n}]\backslash\{c_{2}\}$-distinguishing proper vertex coloring of $G$. Define $C_{1}=f[\bigcup_{n\in\omega}A_{n}]\backslash\{c_{2}\}$. By Claim 4.4(4), $C_{1}\prec f[\bigcup_{n\in\omega}A_{n}]\preceq C$ which contradicts the fact that $\kappa$ is the distinguishing chromatic number of $G$. ∎
4.3. A Lemma on edge colorings
By Claim 4.2, every automorphism fixes the edges $\{t^{\prime\prime},t^{\prime}\}$, $\{t^{\prime},t_{0}\}$, and $\{t_{n},t_{n+1}\}$ for each $n\in\omega$. The following lemma will be applied in subsections 4.4, 4.5, 4.6, and 4.7. Throughout the section, let $B=\{\{t_{n},x\}:n\in\omega,x\in A_{n}\}$.
Lemma 4.6.
If $f$ is a distinguishing proper edge coloring (resp. neighbor-distinguishing edge coloring, total coloring) of $G$, then the following holds:
  1. (1)
    For each $n\in\omega$ and $x,y\in A_{n}$ such that $x\neq y$, $f(\{t_{n},x\})\neq f(\{t_{n},y\})$.
  2. (2)
    $f[B]$ is an infinite, Dedekind-finite set.
  3. (3)
    For each natural number $m$, if $c_{0},...,c_{m}\in f[B]$, then there is an edge coloring $h:B\to f[B]\setminus\{c_{0},...,c_{m}\}$ such that $h(x)=f(x)$ whenever $f(x)\neq c_{i}$.
Proof.
This follows modifying the arguments of Claims 4.3, and 4.4. ∎
4.4. G has no distinguishing chromatic index
Assume that the graph $G$ has a distinguishing chromatic index. Let $f:E_{G}\to C$ be a distinguishing proper edge coloring with $|C|=\kappa$, where $\kappa$ is the distinguishing chromatic index of $G$. By Lemma 4.6, $f[B]$ is an infinite, Dedekind-finite set. Moreover, for some $c_{0},c_{1},c_{2}\in f[B]$, we can define an edge coloring $h:B\to f[B]\setminus\{c_{0},c_{1},c_{2}\}$ as in Lemma 4.6(3). Define,
  1. (1)
    $h(\{t^{\prime\prime},t^{\prime}\})=c_{0}$, $h(\{t^{\prime},t_{0}\})=c_{1}$,
  2. (2)
    $h(\{t_{2n},t_{2n+1}\})=c_{0}$ and $h(\{t_{2n+1},t_{2n+2}\})=c_{1}$ for all $n\in\omega$.
Thus, we obtain a $f[B]\setminus\{c_{2}\}$- distinguishing proper edge coloring $h:E_{G}\to f[B]\setminus\{c_{2}\}$, with $f[B]\setminus\{c_{2}\}\prec f[B]\preceq C$ as $f[B]$ is Dedekind-finite, contradicting the fact that $\kappa$ is the distinguishing chromatic index of $G$.
4.5. G has no neighbor-distinguishing index
Assume that $G$ has a neighbor-distinguishing index. Let $f:E_{G}\to C$ be a neighbor-distinguishing edge coloring with $|C|=\kappa$, where $\kappa$ is the distinguishing chromatic index of $G$. By Lemma 4.6, $f[B]$ is an infinite, Dedekind-finite set and for some $c_{0},c_{1},c_{2},c_{3}\in f[B]$, we can define an edge coloring $h:B\to f[B]\setminus\{c_{0},c_{1},c_{2},c_{3}\}$ as in Lemma 4.6(3). Define,
  1. (1)
    $h(\{t^{\prime\prime},t^{\prime}\})=c_{1}$, $h(\{t^{\prime},t_{0}\})=c_{2}$,
  2. (2)
    $h(\{t_{3n},t_{3n+1}\})=c_{0}$, $h(\{t_{3n+1},t_{3n+2}\})=c_{1}$, and $h(\{t_{3n+2},t_{3n+3}\})=c_{2}$ for all $n\in\omega$.
Thus, we obtain a $f[B]\setminus\{c_{3}\}$-neighbor-distinguishing edge coloring $h:E_{G}\to f[B]\setminus\{c_{3}\}$, with $f[B]\setminus\{c_{3}\}\prec f[B]\preceq C$ as $f[B]$ is Dedekind-finite, contradicting the fact that $\kappa$ is the neighbor-distinguishing index of $G$.
4.6. G has no total chromatic number
Assume that $G$ has a total chromatic number. Let $f:V_{G}\cup E_{G}\to C$ be a total coloring with $|C|=\kappa$, where $\kappa$ is the total chromatic number of $G$. By Lemma 4.6, $f[B]$ is an infinite, Dedekind-finite set and for some $c_{i}\in f[B]$ where $0\leq i\leq 5$, we can define an edge coloring $h:B\to f[B]\setminus\{c_{0},c_{1},c_{2},c_{3},c_{4},c_{5}\}$ as in Lemma 4.6(3). Let,
  1. (1)
    $h(\{t^{\prime\prime},t^{\prime}\})=c_{0}$ and $h(\{t^{\prime},t_{0}\})=c_{1}$.
  2. (2)
    $h(\{t_{2n},t_{2n+1}\})=c_{0}$ and $h(\{t_{2n+1},t_{2n+2}\})=c_{1}$ for all $n\in\omega$.
  3. (3)
    $h(t^{\prime\prime})=c_{2}$ and $h(t^{\prime})=c_{3}$.
  4. (4)
    $h(t_{2n})=c_{2}$ and $h(t_{2n+1})=c_{3}$ for all $n\in\omega$.
  5. (5)
    $h(x)=c_{4}$ for all $x\in\bigcup_{i\in\omega}A_{i}$.
Then, $h:V_{G}\cup E_{G}\rightarrow f[B]\backslash\{c_{5}\}$ is a $f[B]\setminus\{c_{5}\}$-total coloring of $G$. Finally, $f[B]\setminus\{c_{5}\}\prec f[B]\preceq C$ contradicts the fact that $\kappa$ is the total chromatic number of $G$.
4.7. G has no total distinguishing chromatic number
Assume that $G$ has a total distinguishing chromatic number, and let $f:V_{G}\cup E_{G}\to C$ be a properly total coloring with $|C|=\kappa$ such that only the trivial automorphism preserves it, where $\kappa$ is the total distinguishing chromatic number of $G$. We claim that the coloring $h$ in subsection 4.6 is defined in a way so that only the trivial automorphism preserves $h$. In particular, by Claim 4.2, the vertices $t^{\prime},t^{\prime\prime}$, and $t_{m}$ are fixed by every automorphism for each non-negative integer $m$. Since $h$ is a properly total coloring, any neighbor of $t_{m}$ has to be mapped into itself because every edge incident to $t_{m}$ has a distinct color by Lemma 4.6(1). Thus, following the arguments of subsection 4.6, we obtain a contradiction.
5. Odd colorings
Theorem 5.1.
($\mathsf{ZF}$)The following statements are equivalent:
  1. (1)
    Kőnig’s Lemma.
  2. (2)
    Every infinite locally finite connected graph has an odd chromatic number.
Proof.
Implication (1$\Rightarrow$(2) follows from Observation 3.2, and the fact that $\mathsf{AC_{fin}^{\omega}}$ implies every infinite locally finite connected graph is countably infinite.
(2$\Rightarrow$(1) In view of the proof of Theorem 4.1, it suffices to show that the statement mentioned in (2) implies $\mathsf{PAC_{fin}^{\omega}}$. Assume $\mathcal{A}$ and $T$ as in the proof of Theorem 4.1.
[…]
Figure 3. Graph $G_{1}$, an infinite locally finite connected graph.
Consider the following infinite locally finite connected graph $G_{1}=(V_{G_{1}},E_{G_{1}})$ (see Figure 3):
$V_{G_{1}}$ $:=(\bigcup_{n\in\omega}A_{n})\cup T,$
$E_{G_{1}}$ :=$:=\big{\{}\{t_{n},t_{n+1}\}:n\in\omega\big{\}}$$\cup\big{\{}\{t_{n},x\}:n\in\omega,x\in A_{n}\big{\}}$$\cup\big{\{}\{x,y\}:n\in\omega,x,y\in A_{n},x\neq y\big{\}}.$
Claim 5.2.
If $G_{1}$ has a $C$-odd proper vertex coloring $f:V_{G_{1}}\rightarrow C$, then the following holds:
  1. (1)
    $f[\bigcup_{n\in\omega}A_{n}]$ is an infinite, Dedekind-finite set.
  2. (2)
    For each natural number $m$, if $c_{0},...,c_{m}\in f[\bigcup_{n\in\omega}A_{n}]$, then there is a vertex coloring $h:\bigcup_{n\in\omega}A_{n}\to f[\bigcup_{n\in\omega}A_{n}]\setminus\{c_{0},...,c_{m}\}$ such that $h(x)=f(x)$ whenever $f(x)\neq c_{i}$.
Proof.
This follows modifying the arguments of Claims 4.3, and 4.4 (cf. [3, the proof of Theorem 4.2] as well as [12, the proof of Proposition 23]). ∎
Claim 5.3.
$G_{1}$ has no odd chromatic number.
Proof.
Assume that $G_{1}$ has an odd chromatic number. Let $f:V_{G_{1}}\to C$ be a $C$-odd proper vertex coloring with $|C|=\kappa$, where $\kappa$ is the odd chromatic number of $G_{1}$. By Claim 5.2(1), $f[\bigcup_{n\in\omega}A_{n}]$ is an infinite, Dedekind-finite set. Consider the vertex coloring $h:\bigcup_{n\in\omega}A_{n}\to f[\bigcup_{n\in\omega}A_{n}]\setminus\{c_{0},c_{1},c_{2}\}$ for some $c_{0},c_{1},c_{2}\in f[\bigcup_{n\in\omega}A_{n}]$ as in Claim 5.2(2). Let, $h(t_{2n})=c_{0}$, and $h(t_{2n+1})=c_{1}$ for all $n\in\omega$. Then, $h:V_{G_{1}}\rightarrow(f[\bigcup_{n\in\omega}A_{n}]\backslash\{c_{2}\})$ is a $f[\bigcup_{n\in\omega}A_{n}]\backslash\{c_{2}\}$-proper vertex coloring of $G_{1}$. The following arguments show that $h$ is an odd proper vertex coloring:
  1. (1)
    If $t_{i}\in T$, then pick the color $c$ of a random vertex of $A_{i}$, say $a_{i}$. Since $h$ is a proper vertex coloring, $h^{-1}(c)\cap N(t_{i})=\{a_{i}\}$ where $N(v)=\{u\in V_{G_{1}}:\{u,v\}\in E_{G_{1}}\}$ is the open neighborhood of $v$ for any $v\in V_{G_{1}}$.
  2. (2)
    If $x\in A_{m}$ for some $m\in\omega$, then let $c\in\{c_{0},c_{1}\}$ be the color of $t_{m}$. Clearly, $h^{-1}(c)\cap N(x)=\{t_{m}\}$ as neither $c_{0}$ nor $c_{1}$ belongs to $h[\bigcup_{n\in\omega}A_{n}]$ because $h$ is a proper vertex coloring.
We define $C_{1}=f[\bigcup_{n\in\omega}A_{n}]\backslash\{c_{2}\}$. Since $f[\bigcup_{n\in\omega}A_{n}]$ is Dedekind-finite, $C_{1}\prec f[\bigcup_{n\in\omega}A_{n}]\preceq C$. This contradicts the fact that $\kappa$ is the odd chromatic number of $G_{1}$. ∎
6. Brooks type upper bounds fails in ZF
Theorem 6.1.
($\mathsf{ZF}$) The following statements fail:
  1. (1)
    If $G$ is a connected infinite graph where $\Delta(G)$ is finite, then $\chi(G)\leq\Delta(G)$,
  2. (2)
    If $G$ is a connected infinite graph where $\Delta(G)$ is finite, then $\chi^{\prime}(G)\leq\Delta(G)+1$,
  3. (3)
    If $G$ is a connected infinite graph where $\Delta(G)$ is finite, then $D(G)\leq\Delta(G)$,
  4. (4)
    If $G$ is a connected infinite graph where $\Delta(G)\geq 3$ is finite, then $D(G)\leq\Delta(G)-1$,
  5. (5)
    If $G$ is a connected infinite graph where $\Delta(G)\geq 3$ is finite, then $D^{\prime}(G)\leq\Delta(G)-1$,
  6. (6)
    If $G$ is a connected infinite graph where $\Delta(G)$ is finite, then $\chi_{D}(G)\leq 2\Delta(G)-1$,
  7. (7)
    If $G$ is a connected infinite graph where $\Delta(G)$ is finite, then $\chi^{\prime}_{D}(G)\leq\Delta(G)+2$,
  8. (8)
    If $G$ is a connected infinite graph where $\Delta(G)$ is finite, then $\chi^{\prime\prime}(G)\leq\Delta(G)+10^{26}$,
  9. (9)
    If $G$ is a connected infinite graph where $\Delta(G)$ is finite, then $\chi^{\prime\prime}_{D}(G)\leq\Delta(G)+10^{26}+1$,
  10. (10)
    If $G$ is a connected infinite graph where $\Delta(G)$ is finite, then $\chi_{o}(G)\leq 2\Delta(G)$,
  11. (11)
    If $G$ is a connected infinite graph where $\Delta(G)$ is finite, then $D^{\prime\prime}(G)\leq\lceil\sqrt{\Delta(G)}\rceil$.
Proof.
Let $\mathcal{A}=\{A_{n}:n\in\omega\}$ be a countably infinite family of $n$-element sets without a partial choice function for some natural number $n\geq 2$. Consider the graphs $G$ and $G_{1}$ from the proof of Theorems 4.1 and 5.1 (see Figures 2 and 3). Clearly, $\Delta(G)=\Delta(G_{1})=n+2$. By the arguments in the proof of Theorems 4.1 and 5.1, $\chi_{D}(G)$, $\chi^{\prime}_{D}(G)$, $\chi^{\prime\prime}(G)$, $\chi^{\prime\prime}_{D}(G)$, and $\chi_{o}(G_{1})$ do not exist. Let $H_{1}$ be the graph obtained from the graph $G_{1}$ of Figure 3 after deleting the edge set $\{\{x,y\}:n\in\omega,x,y\in A_{n},x\neq y\}$ (see Figure 4). Clearly, $H_{1}$ is a connected infinite graph where $\Delta(H_{1})=n+2$.
[…]
Figure 4. Graph $H_{1}$, an infinite connected graph where $\Delta(H_{1})=n+2$.
By the arguments in the proof of [3, Claims 4.5, 4.6, 4.9, 4.10], $\chi(G_{1})$, $\chi^{\prime}(H_{1}),D(G),$ and $D^{\prime}(G)$ do not exist. Thus, (1)-(10) fails in $\mathsf{ZF}$.
Claim 6.2.
Let $f:G\rightarrow D^{\prime\prime}(G)$ be a total distinguishing coloring of $G$, and $B=\{(\{t_{n},x\},x):n\in\omega,x\in A_{n}\}$. Let $g:B\rightarrow D^{\prime\prime}(G)\times D^{\prime\prime}(G)$ be a function which maps ($(\{t,x\},x)$ to $(f(\{t,x\}),f(x))$. Then the following holds:
  1. (1)
    For each natural number $n$ and for each $x,y\in A_{n}$ such that $x\neq y$, we have $g(\{t_{n},x\},x)\neq g(\{t_{n},y\},y)$ (i.e., $\{f(\{t_{n},x\}),f(x)\}\neq\{f(\{t_{n},y\}),f(y)\}$).
  2. (2)
    $g[B]$ is an infinite set.
Proof.
(1). This is straightforward.
(2). Define, $N_{(c_{1},c_{2})}=\{(\{t_{n},x\},x)\in g^{-1}(c_{1},c_{2}):n\in\omega,x\in A_{n}\}$ for $(c_{1},c_{2})\in ran(g)$. Clearly, $N_{(c_{1},c_{2})}$ is finite, otherwise $\pi_{2}[N_{(c_{1},c_{2})}]$ will generate a partial choice function where $\pi_{2}$ is a function which maps ($(\{t_{n},x\},x)$ to $x$. Since $B=\bigcup_{(c_{1},c_{2})\in g[B]}N_{(c_{1},c_{2})}$ and finite union of finite sets is finite in $\mathsf{ZF}$, if $g[B]$ is finite, then $B$ is finite as well. Thus $g[B]$ is infinite. ∎
We claim that $D^{\prime\prime}(G)$ is infinite (if it exists) and so (11) fails. If $D^{\prime\prime}(G)$ is finite, then $D^{\prime\prime}(G)\times D^{\prime\prime}(G)$ is finite. This contradicts Claim 6.2(2), since $g[B]\subseteq D^{\prime\prime}(G)\times D^{\prime\prime}(G)$. ∎
7. Appendix
In this section, we list the complete definitions of the graph coloring parameters.
Definition 7.1.
A graph $G=(V_{G},E_{G})$ consists of a set $V_{G}$ of vertices and a set $E_{G}\subseteq[V_{G}]^{2}$ of edges (i.e., $E_{G}$ is a subset of the set of all two-element subsets of $V_{G}$). Two vertices $x,y\in V_{G}$ are adjacent vertices if $\{x,y\}\in E_{G}$, and two edges $e,f\in E_{G}$ are adjacent edges if they share a common vertex. The degree of a vertex $v\in V_{G}$, denoted by $deg(v)$, is the number of edges emerging from $v$. We denote by $\Delta(G)$ the maximum degree of $G$. Given a non-negative integer $n$, a path of length $n$ in $G$ is a one-to-one finite sequence $\{x_{i}\}_{0\leq i\leq n}$ of vertices such that for each $i<n$, $\{x_{i},x_{i+1}\}\in E_{G}$; such a path joins $x_{0}$ to $x_{n}$.
  1. (1)
    $G$ is locally finite if every vertex of $G$ has a finite degree.
  2. (2)
    $G$ is connected if any two vertices are joined by a path of finite length.
  3. (3)
    Kőnig’s Lemma states that every infinite locally finite connected graph has a ray.
  4. (4)
    $C_{n}$ is the cycle graph on $n$ vertices for any natural number $n\geq 3$.
  5. (5)
    A proper vertex coloring of $G$ with a color set $C$ is a mapping $f:V_{G}\rightarrow C$ such that for every $\{x,y\}\in E_{G}$, $f(x)\not=f(y)$.
  6. (6)
    A proper edge coloring of $G$ with a color set $C$ is a mapping $f:E_{G}\rightarrow C$ such that for any two adjacent edges $e_{1}$ and $e_{2}$, $f(e_{1})\not=f(e_{2})$.
  7. (7)
    A total coloring of $G$ with a color set $C$ is a mapping $f:V_{G}\cup E_{G}\rightarrow C$ such that for every $\{x,y\}\in E_{G}$, $f(x),f(\{x,y\})$ and $f(y)$ are all different and for any two adjacent edges $e_{1}$ and $e_{2}$, $f(e_{1})\not=f(e_{2})$.
  8. (8)
    A proper vertex coloring $f$ of $G$ is said to be odd if for each non-isolated vertex $x\in V_{G}$ there exists a color $c$ such that $|f^{-1}(c)\cap N(x)|$ is odd where $N(v)=\{u\in V_{G}:\{u,v\}\in E_{G}\}$ is the open neighborhood of $v$ for any $v\in V_{G}$ (cf. [6, Introduction]).
  9. (9)
    Let $|C|=\kappa$. We say $G$ is $\kappa$-proper vertex colorable or $C$-proper vertex colorable if there is a proper vertex coloring $f:V_{G}\rightarrow C$, $G$ is $\kappa$-total colorable or $C$-total colorable if there is a total coloring $f:V_{G}\cup E_{G}\rightarrow C$, $G$ is $\kappa$-odd proper vertex colorable or $C$-odd proper vertex colorable if there is an odd proper vertex coloring $f:V_{G}\rightarrow C$, and $G$ is $\kappa$-proper edge colorable or $C$-proper edge colorable if there is a proper edge coloring $f:E_{G}\rightarrow C$.
    1. (a)
      The least cardinal $\kappa$ for which $G$ is $\kappa$-proper vertex colorable (if it exists) is the chromatic number $(\chi(G))$ of $G$,
    2. (b)
      The least cardinal $\kappa$ for which $G$ is $\kappa$-total colorable (if it exists) is the total chromatic number $(\chi^{\prime\prime}(G))$ of $G$,
    3. (c)
      The least cardinal $\kappa$ for which $G$ is $\kappa$-odd proper vertex colorable (if it exists) is the odd chromatic number $(\chi_{o}(G))$ of $G$,
    4. (d)
      The least cardinal $\kappa$ for which $G$ is $\kappa$-proper edge colorable (if it exists) is the chromatic index $(\chi^{\prime}(G))$ of $G$.
  10. (10)
    For a $\kappa$-proper edge coloring $f$ of $G$, we use $C_{f}(v)$ to denote the set of colors assigned to the edges incident with a vertex $v$. The $\kappa$-proper edge coloring $f$ is called neighbor-distinguishing if $C_{f}(u)\neq C_{f}(v)$ for any pair of adjacent vertices $u$ and $v$. The neighbor-distinguishing index $(\chi^{\prime}_{N}(G))$ of $G$ is the least cardinal $\kappa$ (if it exists) such that $G$ has a $\kappa$-neighbor-distinguishing proper edge coloring (cf. [14, Introduction]).
  11. (11)
    An automorphism of $G$ is a bijection $\phi:V_{G}\rightarrow V_{G}$ such that $\{u,v\}\in E_{G}$ if and only if $\{\phi(u),\phi(v)\}\in E_{G}$. Let $f$ be an assignment of colors to either vertices or edges of $G$. We say that an automorphism $\phi$ of $G$ preserves $f$ if each vertex of $G$ is mapped to a vertex of the same color or each edge of $G$ is mapped to an edge of the same color. We say that $f$ is a distinguishing coloring if the only automorphism that preserves $f$ is the identity. Let $|C|=\kappa$. We say $G$ is $\kappa$-distinguishing vertex colorable or $C$-distinguishing vertex colorable if there is a distinguishing vertex coloring $f:V_{G}\rightarrow C$ and $G$ is $\kappa$-distinguishing edge colorable or $C$-distinguishing edge colorable if there is a distinguishing edge coloring $f:E_{G}\rightarrow C$. We say $G$ is $\kappa$-distinguishing proper vertex colorable or $C$-distinguishing proper vertex colorable if there is a distinguishing proper vertex coloring $f:V_{G}\rightarrow C$ and $G$ is $\kappa$-distinguishing proper edge colorable or $C$-distinguishing proper edge colorable if there is a distinguishing proper edge coloring $f:E_{G}\rightarrow C$ (cf. [15, Introduction]).
    1. (a)
      The least cardinal $\kappa$ for which $G$ is $\kappa$-distinguishing vertex colorable (if it exists) is the distinguishing number $(D(G))$ of $G$,
    2. (b)
      The least cardinal $\kappa$ for which $G$ is $\kappa$-distinguishing edge colorable (if it exists) is the distinguishing index $(D^{\prime}(G))$ of $G$.
    3. (c)
      The least cardinal $\kappa$ for which $G$ is $\kappa$-distinguishing proper vertex colorable (if it exists) is the distinguishing chromatic number $(\chi_{D}(G))$ of $G$.
    4. (d)
      The least cardinal $\kappa$ for which $G$ is $\kappa$-distinguishing proper edge colorable (if it exists) is the distinguishing chromatic index $(\chi^{\prime}_{D}(G))$ of $G$.
  12. (12)
    Let $g$ be an assignment of colors to both vertices or edges of $G$. We say that $g:V_{G}\cup E_{G}\rightarrow C$ is a total distinguishing coloring if $g$ is only preserved by the trivial automorphism. We say that $g$ is a properly total distinguishing coloring if $g$ is total distinguishing and it is a total coloring (see (7)). We say $G$ is $\kappa$-distinguishing total colorable or $C$-distinguishing total colorable if there is a total distinguishing coloring $f:V_{G}\cup E_{G}\rightarrow C$ and $G$ is $\kappa$-properly distinguishing total colorable or $C$-properly distinguishing total colorable if there is a properly total distinguishing coloring $f:V_{G}\cup E_{G}\rightarrow C$ (cf. [15, 18]).
    1. (a)
      The least cardinal $\kappa$ for which $G$ is $\kappa$-distinguishing total colorable (if it exists) is the total distinguishing number ($D^{\prime\prime}(G)$) of $G$.
    2. (b)
      The least cardinal $\kappa$ for which $G$ is $\kappa$-properly distinguishing total colorable (if it exists) is the total distinguishing chromatic number ($\chi^{\prime\prime}_{D}(G)$) of $G$.
  13. (13)
    The automorphism group of $G$, denoted by $Aut(G)$, is the group consisting of automorphisms of $G$ with composition as the operation. Let $\tau$ be a group acting on a set $S$ and let $a\in S$. The orbit of $a$, denoted by $Orb_{\tau}(a)$, is the set $\{\phi(a):\phi\in\tau\}$.